OpenStudy (abbycross167):

Can someone please show me the steps to this question? Please If 3x + 4y = 9, then x equals what

OpenStudy (anonymous):

3x + 4y = 9 Move 4y to the RHS. 3x = 9 - 4y Divide each side by 3. x = (9 - 4y)3

OpenStudy (anonymous):

Sorry. The last line should be: x = (9 - 4y)/3

OpenStudy (abbycross167):

oh okay, then what?

OpenStudy (abbycross167):

@robtobey

OpenStudy (anonymous):

\[x=\frac{1}{3} (9-4 y) \]

OpenStudy (abbycross167):

ok is that all? @robtobey

OpenStudy (anonymous):

Yes. This is a simple two step manipulation problem. Thank you for the medal.

OpenStudy (abbycross167):

oh okay, thank you for helping me!